The three sides of a right-angled triangle are proportional to the numbers 3, 4, and 5. Find the length of each side knowing that the area of the triangle is 24m squared.
Work so far:
1/2bh=A
1/2(3x)(4x)=24m squared

Answer by Alan3354(69443)

The three sides of a right-angled triangle are proportional to the numbers 3, 4, and 5. Find the length of each side knowing that the area of the triangle is 24m squared.
Work so far:
1/2bh=A
1/2(3x)(4x)=24m squared
--------
Continue
6x^2 = 24
x^2 = 4
x = 2
-------
Sides are 6, 8 & 10
--> Area = 6*8/2 = 24